# DirectMultiStep: Direct Route Generation for Multistep Retrosynthesis

### v2

@@ 3.3 Retrosynthetic planning of pharmaceutical compounds
This example, however, does not imply that the model struggles to predict routes with convergent branches. If one defines a convergent route as the one that has at least one node (compound) with at least two children that are not leaves, and repeats the analysis, no significant differences in Top-K accuracies (Fig. S5-S6) or in the number of routes found (Fig. S7-S8) is observed despite significant variation in the number of routes that are convergent among training routes, n1, and n5 (Fig. S3).
@@ 4 Conclusion
Our flagman DMS-Flex (Duo) model, incorporating mixture-of-experts approach, outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set. We demonstrate the quality of predicted routes by testing our models on several FDA-approved drugs. Those predictions also highlight limitations caused by the suboptimal diversity of the current training set, particularly for less common reaction types. Those limitations are partially mitigated by employing mixture-of-experts techniques, making them a promising candidate for further research in retrosynthesis prediction and applications of generative models to chemistry in general.
@@ 5.2 Model Architecture
We present a family of DirectMultiStep (DMS) models with varying parameter counts and architectural details. DMS-Flash and DMS-Deep the classical transformer encoder-decoder architecture [56], whereas DMS-Flex and DMS-Wide incorporate sparsely-gated mixture-of-experts layers [46, 47, 48]. Major parameters choices are given in model card (Table  1). All models use GeLU activations [57].
Intermediate experiments with evaluations on subsets of n1 and n5 with 50 and 500 routes reveal no benefit in increasing the number of hidden dimensions to 512 (Tables S4-S5), which could be rationalized by noticing that 256 is already almost 5x the vocabulary size. Although a feed-forward multiplier of four is more common in NLP models, we find no significant improvement in performance as compared to the feed-forward multiplier of three (Tables S6-S7). In addition, the number of encoder layers does not have to be the same as the number of decoder layers, and models with 6 encoder and 12 decoder layers (trained with SM) performed almost as good as the models with 12 encoder and decoder layers. For MoE models, we find no significant improvement upon increase of number of experts to four, and models trained with two active experts outperform those trained with one.
@@ 5.3 Training
During training, a 10% dropout rate is used in embedding, attention, and feed-forward layers. The AdamW optimizer [58, 2019adamw] is used with a learning rate scheduler that warms up from 0 to $3.5\times 10^{-4}$ for DMS-Flex and DMS-Wide, $3\times 10^{-4}$ for DMS-Flash, and $1\times 10^{-4}$ for DMS-Deep during the first 10% of training steps and undergoes cosine decay to $10\%$ of the maximal learning rate over 20 epochs. The optimal initial learning rate for each model was chosen by starting from 5e-4 and monitoring for gradient explosion (in which case the learning rate was decreased by 0.5e-4). Gradients are clipped to 1.0. In addition, each token in the input to the encoders has a 5% chance of being masked. During hyperparameter optimization, the model was validated on a 5% split from the training routes (which exclude n1 and n5 routes).
@@ 6 Data and Software Availability
YS, HL, and AM contributed equally to this work. YS conceived the idea. YS and HL conducted preliminary model training and testing. AM introduced the mixture of experts (MoE) approach and carried out additional model training and testing. VSB provided guidance for the research. All authors participated in writing and reviewing the manuscript.
@@ 7 Acknowledgement
DMS Model Encoder Decoder Hidden FF Total Active Experts
Van-19 6 6 256 3 1 1 1
Van-60 8 8 512 4 1 1 1
MoE-1/3 6 6 256 3 3 1 1
Figure S7: Distribution of the average number of routes for targets in n1 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock. Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i). A route is considered convergent if any node has at least two children that are not leaves.
Figure S8: Distribution of the average number of routes for targets in n5 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock. Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i). A route is considered convergent if any node has at least two children that are not leaves.
Table S4: Top-$K$ accuracy on route test set-n1 (10 000 routes).
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Van-19 19M 0.44 0.53 0.57 0.59 0.60 0.63
MoE-1/3 38M 0.40 0.47 0.50 0.51 0.52 0.53
Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
Table S5: Top-$K$ accuracy on route test set-n5 (10 000 routes).
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Van-19 19M 0.39 0.47 0.50 0.52 0.53 0.56
MoE-1/3 38M 0.36 0.42 0.45 0.46 0.46 0.48
Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46

### v3

@@ 3.4 Retrosynthetic planning of pharmaceutical compounds
This example, however, does not imply that the model struggles to predict routes with convergent branches. If one defines a convergent route as the one that has at least one node (compound) with at least two children that are not leaves, and repeats the analysis, no significant differences in the number of routes found (Fig. S1-S2) or Top-K accuracies (Fig. S7-S8) is observed despite significant variation in the number of routes that are convergent among training routes, n1, and n5 (Fig. S6).
@@ 4 Conclusion
Our flagman DMS-Flex (Duo) model, incorporating mixture-of-experts approach, outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.5x improvement in Top-1 accuracy on an n1 test set and a 3.9x improvement on the n5 test set. We demonstrate the quality of predicted routes by testing our models on several FDA-approved drugs. Those predictions also highlight limitations caused by the suboptimal diversity of the current training set, particularly for less common reaction types. Those limitations are partially mitigated by employing mixture-of-experts techniques, making them a promising candidate for further research in retrosynthesis prediction and applications of generative models to chemistry in general.
@@ 5.2 Model Architecture
We present a family of DirectMultiStep (DMS) models with varying parameter counts and architectural details. DMS-Flash and DMS-Deep employ the classical transformer encoder-decoder architecture [58], whereas DMS-Flex and DMS-Wide incorporate sparsely-gated mixture-of-experts layers [48, 49, 50]. To create MoE model, one replaces the 2-layer MLP that follows every attention block with an ensemble of 2-layer MLPs (called experts). A preceding router determines (for each input) the $n$-most relevant (active) experts, and passes the input through each of them. Output of each active expert is added with weights determined by the router. Notably, the number of experts active during inference does not have to be equal to the number of experts active during training, although the latter tends to be optimal. Major parameters choices are given in model card (Table  5). All models use GeLU activations [59].
DMS Model Encoder Decoder Hidden FF Total Active Experts
Flash 6 6 256 3 1 1 1
Explorer 6 6 256 3 3 2 2
Explorer XL 8 24 256 3 3 2 2
Flex (Mono) 6 6 256 3 3 2 1
Flex (Duo) 6 6 256 3 3 2 2
Deep 12 36 256 3 1 1 1
Wide 12 12 256 3 3 2 2
Intermediate experiments with evaluations on subsets of n1 and n5 with 50 and 500 routes reveal no benefit in increasing the number of hidden dimensions to 512 (Tables S4-S6), which could be rationalized by noticing that 256 is already almost 5x the vocabulary size. Although a feed-forward multiplier of four is more common in NLP models, we find no significant improvement in performance as compared to the feed-forward multiplier of three (Tables S7-S8). In addition, the number of encoder layers does not have to be the same as the number of decoder layers, and models with 6 encoder and 12 decoder layers (trained with SM) performed almost as good as the models with 12 encoder and decoder layers. For MoE models, we find no significant improvement upon increase of number of experts to four, and models trained with two active experts outperform those trained with one.
@@ 5.3 Training
During training, a 10% dropout rate is used in embedding, attention, and feed-forward layers. The AdamW optimizer [60, 61] is used with a learning rate scheduler that warms up from 0 to $3.5\times 10^{-4}$ for DMS-Flex and DMS-Wide, $3\times 10^{-4}$ for DMS-Flash, and $1\times 10^{-4}$ for DMS-Deep during the first 10% of training steps and undergoes cosine decay to $10\%$ of the maximal learning rate over 20 epochs. The optimal initial learning rate for each model was chosen by starting from 5e-4 and monitoring for gradient explosion (in which case the learning rate was decreased by 0.5e-4). Gradients are clipped to 1.0. In addition, each token in the input to the encoders has a 5% chance of being masked. During hyperparameter optimization, the model was validated on a 5% split from the training routes (which exclude n1 and n5 routes).
@@ 6 Data and Software Availability
YS and AM contributed equally to this work. YS conceived the idea. YS and HL conducted preliminary model training and testing. AM introduced the mixture of experts (MoE) approach. AM and YS carried out additional model training and testing and wrote the manuscript. VSB provided guidance for the research. All authors have given approval to the final version of the paper.
@@ 7 Acknowledgement
Figure S1: Distribution of the average number of routes predicted by DMS Explorer XL for targets in n1 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock. Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i). A route is considered convergent if any node has at least two children that are not leaves.
Figure S2: Distribution of the average number of routes predicted by DMS Explorer XL for targets in n5 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock. Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i). A route is considered convergent if any node has at least two children that are not leaves.
Table S3: Search performance for DirectMultiStep models on USPTO-190 [39]
Methoda Stock Setb Solved Ratec Run Time (s)d
DMS-Explorer-XL eMolecules Screening 33.2% (31.5%) 21.9
MCTS [39] eMolecules Screening 33.7% 370.5
DMS-Flashe,f eMolecules Screening 60.5% (55.2%) 31.3
DMS-Widee eMolecules Screening 63.7% (56.6%) 105.8
Retro* [39] eMolecules Screening 86.8% 156.6
RetroGraph [41] eMolecules Screening 99.5% NA
DMS-Explorer-XL Buyables 27.9% (28.0%) 21.9
Original + MCTS [47] Buyables 46.3% NA
DMS-Flashe,f Buyables 55.3% (51.0%) 31.3
DMS-Widee Buyables 56.8% (51.7%) 105.8
Higherlev + MCTS [47] Buyables 73.7% NA
a All DMS models are run with a beam size of 50 on a single NVIDIA A100 GPU with half-precision floating point inference (fp16).
b eMolecules Screening is from Chen et al. [39] and contains 23.1M screening compounds from eMolecules as of 2019 (2019-11-01). Buyables is from Roh et al. [47] and includes 0.329M buyable building blocks from eMolecules, Sigma-Aldrich, Mcule, ChemBridge Hit2Lead, and WuXi LabNetwork.
c 47 of the targets from USPTO-190 are present as non-leaf nodes in PaRoutes training set. The solved rate on the remaining 143 targets is reported in parentheses.
d Averages over all targets. NA indicates that the run time is not available in the original publication.
e Uses step counts from 2 to 8 (total of 7 DMS model runs).
f SMs not provided for DMS-Flash.
@@
DMS Model Encoder Decoder Hidden FF Total Active Experts
Van-19 6 6 256 3 1 1 1
Van-60 8 8 512 4 1 1 1
MoE-1/3 6 6 256 3 3 1 1
Table S7: Top-$K$ accuracy on route test set-n1 (10 000 routes).
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Van-19 19M 0.44 0.53 0.57 0.59 0.60 0.63
MoE-1/3 38M 0.40 0.47 0.50 0.51 0.52 0.53
Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
Table S8: Top-$K$ accuracy on route test set-n5 (10 000 routes).
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Van-19 19M 0.39 0.47 0.50 0.52 0.53 0.56
MoE-1/3 38M 0.36 0.42 0.45 0.46 0.46 0.48
Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46